OpenStudy (anonymous):

if y = -12 when x = 9, find y when x = -4

OpenStudy (blurbendy):

-12 / 9 = -4 / 3 x / -4 = -4 / 3 Cross multiply and get 3x = 16 x = 16/3

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

find y when x = 10 if y = 8 when x= 20 (using y=kx)

OpenStudy (blurbendy):

8 = 20k k = 8 / 20 k = 2 / 5 y = 2/5(10) y = 4

OpenStudy (anonymous):

Thanks